OpenStudy (barrelracing):

Which of the following statements have the same result? Explain each step in solving each one. f(2) when f(x) = 3x + 2 f−1(3) when f(x) = 2 x minus 7, all over 3 2y + 14 = 4y − 2

OpenStudy (barrelracing):

please help me

OpenStudy (barrelracing):

@freckles

OpenStudy (michaelac959):

do you know how to do these?

OpenStudy (barrelracing):

no

OpenStudy (misty1212):

HI!!

OpenStudy (barrelracing):

hi

OpenStudy (michaelac959):

hi!!!!!!

OpenStudy (misty1212):

\[f(2)\] when \[f(x)=3x+2\] is \[f(2)=3\times 2+2=8\]

OpenStudy (barrelracing):

how did you find that can you show your work so i can better understand it please

OpenStudy (michaelac959):

what grade r u in

OpenStudy (michaelac959):

i did this for algabra 1 in the 8th grade

OpenStudy (misty1212):

\[f^{-1}(3)\] when \[f(x)=\frac{2x-7}{3}\]means solve \[\frac{2x-7}{3}=3\]

OpenStudy (michaelac959):

i can go get my book and explain it

OpenStudy (misty1212):

@Barrelracing are you asking how i found \(f(2)\)?

OpenStudy (barrelracing):

How did you find anyone of the problems? like can you show the work please?

OpenStudy (michaelac959):

what math is this?

OpenStudy (barrelracing):

algebra 2

OpenStudy (misty1212):

there are 3 questions the first one is to find \(f(2)\) for \(f(x)=3x+2\)

OpenStudy (michaelac959):

ok gimme a sec to find my algebra 2 book

OpenStudy (misty1212):

do do that, where you see an \(x\) replace it by a \(2\)

OpenStudy (misty1212):

since \[f(x)=3x+2\] you have \[f(2)=3(2)+2=6+2=8\]

OpenStudy (michaelac959):

ok i got it

OpenStudy (michaelac959):

here is the page

OpenStudy (michaelac959):

it explains how to do it

OpenStudy (barrelracing):

makes since what about the other 2?

OpenStudy (barrelracing):

Misty i really need your help. i am not cheating for anything this is just an example in my lesson i didnt understand @misty1212

OpenStudy (barrelracing):

for anyone who needs the answer f(x) = 3x + 2 f(2) = 3(2) + 2 6+2 8 2(3) - 7/3 6 - 7/3 -1/3 -3 2y + 14 = 4y - 2 16 = 2y 8 = y